# On the adjoint Selmer groups of semi-stable elliptic curves and Flach’s zeta elements

### v4

@@ Abstract.
We explicitly construct the rank one primitive Stark (equivalently, Kolyvagin) system extending a constant multiple of Flach’s zeta elements for semi-stable elliptic curves. As its arithmetic applications, we obtain the equivalence between a specific behavior of the Stark system and the minimal modularity lifting theorem, and we also discuss the cyclicity of the adjoint Selmer groups. Our Stark system construction yields a more refined interpretation of the collection of Flach’s zeta elements than the “geometric Euler system” approach due to Flach, Wiles, Mazur, and Weston.
@@ Introduction
As explained in Introduction of his celebrated work [Wil95], A. Wiles first tried to prove the semi-stable Shimura–Taniyama conjecture and Fermat’s last theorem via the construction of an Euler system for symmetric squares of semi-stable elliptic curves extending the cohomology classes constructed by M. Flach [Fla92], which we call Flach’s zeta elements in this article. However, the construction of the Euler system was incomplete, and it was regarded as a gap.
@@
The main goals of this article are to investigate the behavior of the collection of Flach’s zeta elements beyond the scope of this geometric Euler system strategy and to deduce certain structural results for adjoint Selmer groups. However, it is well-known that the construction question of an Euler system extending Flach’s zeta element is extremely difficult (see [Rub00, §$\S$3.6], [Wes01, pp. 359–360], [MR04, Rem. 6.3.4], and [LZ19, Intro.] for example).
@@
Since there is a certain constant multiple subtlety in our Stark system approach, our result does not give a new proof of the semi-stable Shimura–Taniyama conjecture or Fermat’s last theorem yet. Also, we do not know whether our Stark system actually comes from an Euler system [^2] We are not sure whether it is possible to recover an Euler system from the Kolyvagin system associated to the Euler system.. Indeed, we prove that the minimal modularity lifting theorem is equivalent to showing that the constant multiple is a $p$-adic unit.
@@
The cyclicity question has long history in Iwasawa theory. This is just a quick summary of the preface of Hida’s book [Hid22a] and we strongly recommend the reader to read it.
K. Iwasawa himself was interested in the structure of various classical Iwasawa modules including the cyclicity over the Iwasawa algebra. In [Iwa69], he obtained the cyclicity of certain classical Iwasawa modules over the Iwasawa algebra (as well as the main conjecture) under the Kummer–Vandiver conjecture. Even without assuming the Kummer–Vandiver conjecture, he proposed several interesting questions on the cyclicity around 1980’s [Iwa88, Iwa14].
@@
Recent years, some groups of mathematicians have developed new approaches towards the construction of symmetric square and adjoint Euler systems from different sources [LZ19, Urb21, LZ, SS]. We provide brief remarks on their constructions and highlight the novelty of our approach in contrast to theirs, as per the request to address this aspect.
In [LZ19], an Euler system is constructed from Beilinson–Flach’s elements and bounds Selmer groups of the symmetric square representations of an ordinary modular form twisted by a non-trivial Dirichlet character. This twist is inevitable to have the correct Euler system relation. In [Urb21], an Euler system for adjoint representations of ordinary Hilbert modular forms is constructed directly from congruence modules, but the $R=\mathbb{T}$ theorem is an ingredient of the construction. Also, the corresponding explicit reciprocity law depends on a conjecture on the Fitting ideals of some equivariant congruence modules for abelian base change. In [LZ], an Euler system for adjoint representations of ordinary modular forms is constructed from Asai–Flach classes and is used to prove one-sided divisibility of the main conjecture for the symmetric square Selmer group over the cyclotomic $\mathbb{Z}_{p}$-extension of $\mathbb{Q}$ for all but finitely many height one prime ideals.
In [SS], an Euler system for the symmetric square representation of an ordinary modular form twisted by an odd Dirichlet character is initially constructed by utilizing the pull-back of Eisenstein series on $\mathrm{SO}(3,2)$. Then by using Hida theory (among others), an Euler system for the symmetric square representation (with no extra twist) is obtained. As its consequence, the $R=\mathbb{T}$ theorem follows. The $p$-ordinary condition is assumed in all the work mentioned above.
@@ 1. Statement of the main result
Let $E$ be a semi-stable modular [^3] We do not use the semi-stable modularity theorem of Wiles and Taylor–Wiles [Wil95, TW95] and its generalizations in this article. elliptic curve over $\mathbb{Q}$ of conductor $N$ and $p\geq 5$ a good reduction prime for $E$. Denote by $T_{p}E$ the $p$-adic Tate module for $E$. Assume that $\overline{\rho}:G_{\mathbb{Q}}=\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\mathrm{Aut}_{\mathbb{F}_{p}}(E[p])$ is irreducible and $N=N(\overline{\rho})$ where $N(\overline{\rho})$ is the conductor of $\overline{\rho}$.
@@
$\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}$
for $\mathrm{Sym}^{2}(T_{p}E)$ extending a constant multiple of Flach’s zeta element $c(\ell)$ (recalled in Theorem 2.2) which determines the structure of the $\ell$-strict (trace zero) adjoint Bloch–Kato Selmer group
@@
In particular, the constant multiple is independent of $\ell$. We call $\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}$ the Flach–Stark system.
@@
The explicit construction of $\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}$ is given in §5. ∎
In the rest of this section, we fix a prime $\ell$ satisfying the conditions in Theorem 1.1. Let $\boldsymbol{\kappa}^{\mathrm{Flach},\ell}$ be the (non-trivial and primitive) Flach–Kolyvagin system corresponding to $\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}$ under the isomorphism between the modules of Stark systems and Kolyvagin systems under our setting [MR16, Rem. 11.5 and Thm. 12.4]. Indeed, the constant multiple in Theorem 1.1 is designed to deduce the following formula for the exact bound of the genuine Bloch–Kato Selmer group in terms of $\boldsymbol{\kappa}^{\mathrm{Flach},\ell}$.
@@
$\mathrm{loc}^{s}_{\ell}(\boldsymbol{\kappa}^{\mathrm{Flach},\ell})=\left\{\mathrm{loc}^{s}_{\ell}(\kappa^{\mathrm{Flach},\ell}_{n}):n\in\mathcal{N}^{\mathrm{Flach},(\ell)}_{1}\right\}$
@@
$\mathrm{ord}_{p}(\mathrm{loc}^{s}_{\ell}(\kappa^{\mathrm{Flach},\ell}_{1}))=\mathrm{length}_{\mathbb{Z}_{p}}(\mathrm{Sel}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}])))$
where $\mathrm{Sel}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}]))$ is the adjoint Bloch–Kato Selmer group and $\mathrm{ord}_{p}(\mathrm{loc}^{s}_{\ell}(\kappa^{\mathrm{Flach},\ell}_{1}))$ is the $p$-divisibility index of $\mathrm{loc}^{s}_{\ell}(\kappa^{\mathrm{Flach},\ell}_{1})$ in $\mathrm{H}^{1}_{/f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(T_{p}E))$.
@@
It is known that the primitivity of a (rank one) Kolyvagin system is closely related to showing the corresponding main conjecture [MR04, KKS20, Sak22, Kim]. This principle would apply to our setting as follows if such a formulation is valid. Let $\mathbb{Q}_{\infty}$ be the cyclotomic $\mathbb{Z}_{p}$-extension of $\mathbb{Q}$ and $\Lambda$ the cyclotomic Iwasawa algebra. If the $\Lambda$-adic deformation $\boldsymbol{\kappa}^{\mathrm{Flach},\ell,\infty}$ of $\boldsymbol{\kappa}^{\mathrm{Flach},\ell}$ exists and $\kappa^{\mathrm{Flach},\ell,\infty}_{1}$ is non-zero in $\widehat{\mathrm{Sel}_{\ell\textrm{-}\mathrm{rel}}}(\mathbb{Q}_{\infty},\mathrm{Sym}^{2}(T_{p}E))$, then the hypothetical main conjecture
``$\mathrm{char}_{\Lambda}\left(\widehat{\mathrm{Sel}_{\ell\textrm{-}\mathrm{rel}}}(\mathbb{Q}_{\infty},\mathrm{Sym}^{2}(T_{p}E))/\Lambda\kappa^{\mathrm{Flach},\ell,\infty}_{1}\right)=\mathrm{char}_{\Lambda}\left(\mathrm{Sel}_{\ell\textrm{-}\mathrm{str}}(\mathbb{Q}_{\infty},\mathrm{ad}^{0}(E[p^{\infty}]))^{\vee}\right)"$
@@
Denote by $\mathrm{deg}(\phi)$ the minimal modular degree of $E$. We obtain the following connection between Flach’s zeta elements and the minimal modularity lifting theorem from Theorem 1.1.
@@
See §$\S$6.2. ∎
Of course, all the statements in Corollary 1.4 are true thanks to the work of Wiles and Taylor–Wiles [Wil95, TW95]. One remarkable feature of Corollary 1.4 is that the modularity lifting theorem can be deduced from a certain property regarding Flach’s zeta elements (finally!).
@@
From the viewpoint of Corollary 1.4, we anticipate that our construction of the Stark system will contribute directly to the revival of Wiles’ abandoned “geometric Euler system” approach to the semi-stable Shimura–Taniyama conjecture in a future.
Although we do not know whether or how $\mathrm{loc}^{s}_{\ell}(\boldsymbol{\kappa}^{\mathrm{Flach},\ell})$ determines the structure of $\mathrm{Sel}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}]))$ in general yet [^5] In [Kim], the self-duality of Galois representations is used in an essential way. We are informed that R. Sakamoto is working on the general structure theorem with rank zero Stark/Kolyvagin systems, and it seems to be closely related to our approach., we still have the following structural application under the $p$-indivisibility assumption on $c(\ell)$.
@@
Suppose that any (equivalent) statement of Corollary 1.4 holds. If we further assume that $c(\ell)\in\mathrm{Sel}_{\ell\textrm{-}\mathrm{rel}}(\mathbb{Q},\mathrm{Sym}^{2}(T_{p}E))$ is not divisible by $p$, then the adjoint Selmer group is cyclic, i.e.
@@
See §$\S$6.3. ∎
This cyclicity can be understood as the structural refinement of the Bloch–Kato conjecture (cf. [Kim, Kim24]). We make the following question (not conjecture).
@@
The higher weight generalization of Theorem 1.1 and its interpretation in terms of rank zero Stark/Kolyvagin systems will be investigated in a joint work with Ryotaro Sakamoto.
@@ 2. Flach’s zeta elements and the explicit reciprocity law
Let $k\geq 1$ be an integer. A prime $\ell$ is said to be a $k$-Flach prime if $(\ell,Np)=1$, $\ell\equiv 1\pmod{p^{k}}$, and $a_{\ell}(E)\not\equiv\pm 2\pmod{p}$. This condition is equivalent to that $(\ell,Np)=1$ and the arithmetic Frobenius at $\ell$ acts on $E[p^{k}]$ as $\tau$ chosen in Choice 3.4 below. We call 1-Flach primes by Flach primes for convenience. Denote by $\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}$ the set of all $k$-Flach primes except $\ell$ for fixed $(E,p)$ and by $\mathcal{N}^{\mathrm{Flach},(\ell)}_{k}$ the set of square-free products of primes in $\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}$.

### v6

@@ Abstract.
We explicitly construct the rank one primitive Stark (equivalently, Kolyvagin) system extending a constant multiple of Flach’s zeta elements for semi-stable elliptic curves. As its arithmetic applications, we obtain the equivalence between the $p$-indivisibility of the constant multiple and the minimal modularity lifting theorem, and we also discuss the cyclicity of the adjoint Selmer groups. In particular, we give an affirmative answer to a question of Mazur and Rubin. Our Stark system construction yields a more refined interpretation of the collection of Flach’s zeta elements than the “geometric Euler system” approach due to Flach, Wiles, Mazur, and Weston.
@@ Introduction
As explained in the Introduction of his celebrated work [Wil95], A. Wiles first tried to prove the semi-stable Shimura–Taniyama conjecture and Fermat’s last theorem via the construction of an Euler system for symmetric squares of semi-stable elliptic curves extending the cohomology classes constructed by M. Flach [Fla92], which we call Flach’s zeta elements in this article. However, the construction of the Euler system was incomplete, and it was regarded as a gap.
@@
The main goals of this article are to investigate the behavior of the collection of Flach’s zeta elements beyond the scope of this geometric Euler system strategy and to deduce certain structural results for adjoint Selmer groups. However, it is well-known that the construction question of an Euler system extending Flach’s zeta element is extremely difficult (see [Rub00, $\S$3.6], [Wes01, pp. 359–360], [MR04, Rem. 6.3.4], and [LZ19, Intro.] for example).
@@
Since there is a certain constant multiple subtlety in our Stark system approach, our result does not give a new proof of the semi-stable Shimura–Taniyama conjecture or Fermat’s last theorem yet. Indeed, we prove that the minimal modularity lifting theorem is equivalent to showing that the constant multiple is a $p$-adic unit. Also, we do not know whether our Stark system actually comes from an Euler system [^2] We are not sure whether it is possible to recover an Euler system from the Kolyvagin system associated to the Euler system..
@@
The cyclicity question has a long history in Iwasawa theory. This is just a quick summary of the preface of Hida’s book [Hid22a] and we strongly recommend the reader to read it.
K. Iwasawa himself was interested in the structure of various classical Iwasawa modules including the cyclicity over the Iwasawa algebra. In [Iwa69], he obtained the cyclicity of certain classical Iwasawa modules over the Iwasawa algebra (as well as the main conjecture) under the Kummer–Vandiver conjecture. Even without assuming the Kummer–Vandiver conjecture, he proposed several interesting questions on the cyclicity around 1980 [Iwa88, Iwa14].
@@
In recent years, some groups of mathematicians have developed new approaches towards the construction of symmetric square and adjoint Euler systems from different sources [LZ19, Urb21, LZ, SS]. We provide brief remarks on their constructions and explain why they are all independent of ours.
In [LZ19], the Beilinson–Flach Euler system for the Rankin–Selberg product of the same $p$-ordinary modular form is studied and an is constructed from Beilinson–Flach’s elements and bounds Selmer groups of the symmetric square representations of an ordinary modular form twisted by a non-trivial Dirichlet character. This twist is inevitable to have the correct Euler system relation. In [Urb21], an Euler system for adjoint representations of ordinary Hilbert modular forms is constructed directly from congruence modules, but the $R=\mathbb{T}$ theorem is an ingredient of the construction. Also, the corresponding explicit reciprocity law depends on a conjecture on the Fitting ideals of some equivariant congruence modules for abelian base change. In [LZ], an Euler system for adjoint representations of ordinary modular forms is constructed from Asai–Flach classes and is used to prove one-sided divisibility of the main conjecture for the symmetric square Selmer group over the cyclotomic $\mathbb{Z}_{p}$-extension of $\mathbb{Q}$ up to a power of $p$. In [SS], an Euler system for the symmetric square representation of an ordinary modular form twisted by an odd Dirichlet character is initially constructed by utilizing the pull-back of Eisenstein series on $\mathrm{SO}(3,2)$. Then by using Hida theory (among others), an Euler system for the symmetric square representation (with no extra twist) is obtained. As its consequence, the $R=\mathbb{T}$ theorem follows. The $p$-ordinary condition is assumed in all the work mentioned above.
@@ 1. Statement of the main result
Let $E$ be a semi-stable modular [^3] We do not use the semi-stable modularity theorem of Wiles and Taylor–Wiles [Wil95, TW95] and its generalizations in this article. elliptic curve over $\mathbb{Q}$ of conductor $N$ and $p\geq 5$ a good reduction prime for $E$. Denote by $T_{p}E$ the $p$-adic Tate module for $E$. Assume that $\overline{\rho}:G_{\mathbb{Q}}=\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\mathrm{Aut}_{\mathbb{F}_{p}}(E[p])$ is irreducible and $N=N(\overline{\rho})$ where $N(\overline{\rho})$ is the conductor of $\overline{\rho}$. For $k\geq 1$, let $\mathrm{ad}^{0}(E[p^{k}])$ is the (trace zero) adjoint representation associated to $E[p^{k}]$.
[…]
$\epsilon^{\ell}_{n}\in\bigwedge^{1+\nu(n)}\mathrm{Sel}_{\mathcal{F}^{\ell n}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\otimes\bigwedge^{\nu(n)}\bigoplus_{q|n}\mathrm{Hom}\left(\frac{\mathrm{H}^{1}(\mathbb{Q}_{q},\mathrm{Sym}^{2}(E[p^{k}]))}{\mathrm{H}^{1}_{f}(\mathbb{Q}_{q},\mathrm{Sym}^{2}(E[p^{k}]))},\mathbb{Z}/p^{k}\mathbb{Z}\right)$
where $n$ is a square-free product of the primes in $\mathcal{P}^{\mathrm{Flach},(\ell)}_{k},$ and $\nu(n)$ is the number of prime divisors of $n$, $\mathcal{F}^{\ell n}_{\mathrm{BK}}$ is the $\ell n$-relaxed Bloch–Kato Selmer structure, and $\mathrm{H}^{1}_{f}$ means the finite part of the local cohomology group at $q$ dividing $n$. A Stark system for $\mathrm{Sym}^{2}(T_{p}E)$ is defined by taking the inverse limit.
@@
$\boldsymbol{\epsilon}^{\mathrm{Flach}^{\prime},\ell}$
for $\mathrm{Sym}^{2}(T_{p}E)$ extending a constant multiple of Flach’s zeta element $c(\ell)$ (recalled in Theorem 2.2) which determines the structure of the $\ell$-strict adjoint Bloch–Kato Selmer group
@@
In particular, the constant multiple is independent of $\ell$.
We call $\boldsymbol{\epsilon}^{\mathrm{Flach}^{\prime},\ell}$ the primitive normalization of the Flach–Stark system $\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}$.
@@
We explicitly construct $\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}$ from Flach’s zeta elements and define $\boldsymbol{\epsilon}^{\mathrm{Flach}^{\prime},\ell}$ by multiplying the constant multiple in §5. ∎
In the rest of this section, we fix a prime $\ell$ satisfying the conditions in Theorem 1.1. Let $\boldsymbol{\kappa}^{\mathrm{Flach}^{\prime},\ell}$ be the the primitive normalization of the (non-trivial) Flach–Kolyvagin system $\boldsymbol{\kappa}^{\mathrm{Flach},\ell}$ corresponding to $\boldsymbol{\epsilon}^{\mathrm{Flach}^{\prime},\ell}$ under the isomorphism between the modules of Stark systems and Kolyvagin systems under our setting [MR16, Rem. 11.5 and Thm. 12.4]. Indeed, the constant multiple in Theorem 1.1 is designed to deduce the following formula for the exact bound of the genuine Bloch–Kato Selmer group in terms of $\boldsymbol{\kappa}^{\mathrm{Flach}^{\prime},\ell}$. Let
$\mathrm{loc}^{s}_{\ell}:\mathrm{H}^{1}(\mathbb{Q},\mathrm{Sym}^{2}(T_{p}E))\to\mathrm{H}^{1}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(T_{p}E))\to\mathrm{H}^{1}_{/f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(T_{p}E))$
be the singular quotient of the localization map at $\ell$ where the first map is the localization at $\ell$ and the second map is the natural quotient, and $\mathrm{H}^{1}_{/f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(T_{p}E))=\frac{\mathrm{H}^{1}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(T_{p}E))}{\mathrm{H}^{1}_{f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(T_{p}E))}$.
@@
$\mathrm{loc}^{s}_{\ell}(\boldsymbol{\kappa}^{\mathrm{Flach}^{\prime},\ell})=\left\{\mathrm{loc}^{s}_{\ell}(\kappa^{\mathrm{Flach},\ell}_{n}):n\in\mathcal{N}^{\mathrm{Flach},(\ell)}_{1}\right\}$
@@
$\mathrm{ord}_{p}(\mathrm{loc}^{s}_{\ell}(\kappa^{\mathrm{Flach}^{\prime},\ell}_{1}))=\mathrm{length}_{\mathbb{Z}_{p}}(\mathrm{Sel}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}])))$
[…]
@@
It is known that the primitivity of a (rank one) Kolyvagin system is closely related to showing the corresponding main conjecture [MR04, KKS20, Sak22, Kim25]. This principle would apply to our setting as follows if such a formulation is valid. Let $\mathbb{Q}_{\infty}$ be the cyclotomic $\mathbb{Z}_{p}$-extension of $\mathbb{Q}$ and $\Lambda$ the cyclotomic Iwasawa algebra. If the $\Lambda$-adic deformation $\boldsymbol{\kappa}^{\mathrm{Flach}^{\prime},\ell,\infty}$ of $\boldsymbol{\kappa}^{\mathrm{Flach}^{\prime},\ell}$ exists and $\kappa^{\mathrm{Flach},\ell,\infty}_{1}$ is non-zero in $\widehat{\mathrm{Sel}_{\ell\textrm{-}\mathrm{rel}}}(\mathbb{Q}_{\infty},\mathrm{Sym}^{2}(T_{p}E))$, then the hypothetical main conjecture
$\mathrm{char}_{\Lambda}\left(\widehat{\mathrm{Sel}_{\ell\textrm{-}\mathrm{rel}}}(\mathbb{Q}_{\infty},\mathrm{Sym}^{2}(T_{p}E))/\Lambda\kappa^{\mathrm{Flach}^{\prime},\ell,\infty}_{1}\right)=\mathrm{char}_{\Lambda}\left(\mathrm{Sel}_{\ell\textrm{-}\mathrm{str}}(\mathbb{Q}_{\infty},\mathrm{ad}^{0}(E[p^{\infty}]))^{\vee}\right)$"
@@
Denote by $\mathrm{deg}(\phi)$ the minimal modular degree of $E$ where $\phi:X_{0}(N)\to E$ is the modular parametrization. We obtain the following connection between Flach’s zeta elements and the minimal modularity lifting theorem from Theorem 1.1.
@@
See $\S$6.2. ∎
Of course, all the statements in Corollary 1.4 are true thanks to the work of Wiles and Taylor–Wiles [Wil95, TW95]. One remarkable feature of Corollary 1.4 is that the modularity lifting theorem can be deduced from a certain property regarding Flach’s zeta elements (finally!). More precisely, the final step in deducing the modularity lifting theorem from our Stark system argument is to control the constant multiple.
@@
In this sense, Corollary 1.4 illustrates both the possibility and the limitation of the approach towards modularity lifting theorem based on Flach’s zeta elements.
Although we do not know whether or how $\mathrm{loc}^{s}_{\ell}(\boldsymbol{\kappa}^{\mathrm{Flach}^{\prime},\ell})$ determines the structure of $\mathrm{Sel}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}]))$ in general yet [^5] In [Kim25], the self-duality of Galois representations is used in an essential way. We are informed that R. Sakamoto is working on the general structure theorem with rank zero Stark/Kolyvagin systems, and it seems to be closely related to our approach., we still have the following structural application under the $p$-indivisibility assumption on $c(\ell)$.
@@
Suppose that any (equivalent) statement of Corollary 1.4 holds. If we further assume that $c(\ell)\in\mathrm{Sel}_{\ell\textrm{-}\mathrm{rel}}(\mathbb{Q},\mathrm{Sym}^{2}(T_{p}E))$ is not divisible by $p$ for one $\ell$, then the adjoint Selmer group is cyclic, i.e.
@@
See $\S$6.3. ∎
This cyclicity can be understood as the structural refinement of the Bloch–Kato conjecture (cf. [Kim24, Kim25]). We make the following question (not conjecture).
@@
The higher weight generalization of Theorem 1.1 and its interpretation in terms of rank zero Stark/Kolyvagin systems is being investigated in a joint work in progress with Ryotaro Sakamoto.
@@ 2. Flach’s zeta elements and the explicit reciprocity law
Let $k\geq 1$ be an integer. A prime $\ell$ is said to be a $k$-Flach prime if $(\ell,Np)=1$, $\ell\equiv 1\pmod{p^{k}}$, and $a_{\ell}(E)\not\equiv\pm 2\pmod{p}$. This condition is equivalent to that $(\ell,Np)=1$ and the arithmetic Frobenius at $\ell$ acts on $E[p^{k}]$ as $\tau$ chosen in Choice 3.4 (with fixed $\alpha$) below. We call 1-Flach primes by Flach primes for convenience. Denote by $\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}$ the set of all $k$-Flach primes except $\ell$ for fixed $(E,p)$ and by $\mathcal{N}^{\mathrm{Flach},(\ell)}_{k}$ the set of square-free products of primes in $\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}$.